Write the equation of the line passing through (32,50) with a slope of -11/16 in slope intercept form. Must show all algebraic work. Must show the original substitution of number into the equation and final answer written in slope intercept form with simplified numbers.
Answer by Alan3354(69443)

Write the equation of the line passing through (32,50) with a slope of -11/16 in slope intercept form
------------
Given a point and the slope:
y- y1 = m*(x - x1) where m = slope, and (x1,y1) is the given point.
Rewrite in any form you like.